Chcl2cooh is oxidised to co2,h2o, and cl2 by 600meq of an oxidiser. Same amount of acid can neutralize how many moles of ammonia to form ammonium dichloroacetate..... ........... ANS : 0.1

Grade:12th pass

1 Answers

Vikas TU
5 years ago
Step1. First, write the half-reaction for the oxidation of dichloroaceticacid
CHCl2COOH+2H2O→2CO2+Cl2+6H++6e-
 
Step 2. Calculate the amount of dichloroacetic acid
Six electrons are transferred in the half-reaction, so
1 eq=1/6mol
600 meq = 600×10-3 eq×1/6mol =0.1mol
 
Step 3.mole ratio 1:1 according to the equation below:
CHCl2COOH+NH3→CHCl2COO-lNH+4
Thus moles of acid = moles of ammonium= 0.1 mole.